# Error Analysis Prompting Enables Human-Like Translation Evaluation in Large Language Models

### v3

@@ Abstract
Qingyu ${}^{\diamondsuit,\Re}$,    Baopu ${}^{\flat,\Re}$,  Liang ${}^{\Re}$,  Kanjian ${}^{\diamondsuit}$,  Tom ${}^{\heartsuit}$,  Dacheng ${}^{\Re}$  ${}^{\diamondsuit}$outheast University  ${}^{\Re}$ Explore Academy, JD.com Inc.  ${}^{\flat}$anjing University  ${}^{\heartsuit}$icrosoft [Uncaptioned image] luqingyu@seu.edu.cn, qiubaopu@smail.nju.edu.cn, liangding.liam@gmail.com, tomkocmi@microsoft.com [Uncaptioned image] https://github.com/Coldmist-Lu/ErrorAnalysis_Prompt
@@ 1 Introduction
Large language models (LLMs), especially Generative Pre-trained Transformer (GPT) models Radford et al. (2019); Brown et al. (2020) such as ChatGPT (Ouyang et al., 2022; Achiam et al., 2023), have shown remarkable performance in various natural language processing (NLP) tasks Qin et al. (2023); Zhong et al. (2023). LLMs are capable of integrating multiple NLP tasks and can generate detailed and comprehensive responses to human inquiries. Additionally, they can respond appropriately to follow-up questions and maintain sensitivity throughout several turns of conversation.
@@ 2.1 Translation Evaluation Metric
To verify the reliability of MT metrics, Multi-dimensional Quality Metric (MQM) has been adopted recently in WMT as a high-quality human evaluation strategy (Freitag et al., 2021). It asks human experts to annotate the errors in the hypothesis and categorize them into "Major" and "Minor" indicating their severity. A detailed example of MQM annotation is presented in Appendix A.
@@ 2.3 Error Analysis Prompting
In this work, we combine the CoT and EA paradigms, introducing a novel prompting strategy called Error Analysis Prompting (EAPrompt). As shown in Figure 1, EAPrompt divides the scoring process into two stages: First, the LLM is instructed to identify major and minor errors in the translation ("Instruction: Identify Errors"). Subsequently, the number of these two types of errors is counted ("Instruction: Count Errors"). Distinguished from GEMBA prompting, EAPrompt emulates the evaluation process of MQM and produces more explanable and reliable automatic evaluations.
@@ 2.4 Post-processing of LLM responses
WMT22 En-De 2037 17
conversational, e-commerce, news, social
En-Ru 2037 17
conversational, e-commerce, news, social
Zh-En 1875 20
conversational, e-commerce, news, social
@@ 3.1 Experiment Setup
We utilize the test set from the WMT22 shared tasks (Freitag et al., 2022) in English-German (En-De), English-Russian (En-Ru), and Chinese-English (Zh-En) across 4 different domains - conversational, e-commerce, news, and social. Table 1 provides statistics about our test set.
@@ 3.2 Baselines and Large Language Models
For prioprietary models, we use the OpenAI API to experiment with GPT-3.5-Turbo  [^2] We use the 0613 OpenAI model versions.. We also experiment with a human-aligned Llama2-70B series model Touvron et al. (2023b) fine-tuned with multilingual translation data, noted as "Llama2-70b-Chat" in experimental results. We also use a high-quality sparse mixture-of-experts model, Mixtral-8x7b (Jiang et al., 2024). We use a state-of-the-art checkpoint Mixtral-8x7b-Instruct which has been optimised through supervised fine-tuning and direct preference optimisation to follow instructions.
@@ 3.4 Experimental Results
EAPrompt 89.4 (+2.5) 55.7 (+1.0) 53.4 (+3.4) 48.8 (+1.2)
Table 2: The performance of metrics using pairwise accuracy (%) at the system level and pairwise accuracy with tie calibration (%) at the segment level. All results are compared with human-annotated MQM scores. The best results among the same model are highlighted in bold. The best results among all metrics are underlined.
@@ 4.4 Case Study
We discuss potential issues encountered by LLMs and their corresponding solutions in Appendix E, including invalid responses, input order bias, etc. We aim to provide insights that should be considered when utilizing LLMs as translation evaluators.
@@ 5 Related Work
LLMs refers to language models with hundreds of billion of parameters which are trained on massive textual data (Chang et al., 2024; Zhao et al., 2023). Since the emergence of ChatGPT, LLMs have shown its remarkable proficiency across various NLP tasks (Achiam et al., 2023; Touvron et al., 2023b). A prevalent application of LLMs is harnessing them as evaluators for assessing the performance of Chatbots (Zheng et al., 2023). Recent studies also show LLM’s efficacy in evaluating NLG tasks like summarization and dialog generation through multi-step prompting (Liu et al., 2023). GEMBA (Kocmi and Federmann, 2023b) is the pioneering effort in utilizing LLMs as translation evaluators via a zero-shot prompting approach with GPT models. In this work, EAPrompt innovatively combines error analysis (Lu et al., 2023) and chain-of-thought (Wei et al., 2022) to prompt LLMs for achieving human-like translation evaluation.
@@ 6 Conclusion
In this paper, we explore the potential of LLMs as a metric for evaluating translations. We design a novel one-shot prompting strategy EAPrompt based on chain-of-thought and error analysis, and show that this strategy significantly improves the evaluation performance on both the system and segment levels. We compare different EAPrompt variants and ultimately opt for a 2-step prompting approach with itemized error demonstrations. Further analysis confirms EAPrompt’s proficiency in error identification and its alignment with the commonly accepted human evaluations MQM.
In future work, we would like to experiment with a broader range of LLMs Barrault et al. (2019); Anastasopoulos et al. (2021); Kocmi et al. (2022); Zan et al. (2022), to make our conclusion more convincing. Lastly, it will be interesting to test the capabilities of LLMs for other MT-related tasks, such as grammatical error correction and automatic post-editing Wu et al. (2023); Vidal et al. (2022).
@@ Appendix A Description of MQM
Multidimensional Quality Metric (MQM) is a human evaluation framework commonly used in WMT metrics shared tasks as the golden standard (Freitag et al., 2021, 2023). In this paper, EAPrompt emulates MQM to identify major and minor errors, providing insightful explanations for the translation. Table 6 illustrates an example in detail annotated through MQM framework.
@@ Appendix D Counting Errors using Regular Expressions Matching
Appendix E Case Study
@@
E.1 Potential instability in the responses without temperature control
@@
E.2 Input order bias when evaluating multiple translations simultaneously
@@
E.3 LLMs may generate invalid answers for all prompting strategies
Issue: We observe that in certain cases, LLMs may not function as translation evaluators that may producing invalid answers with textual explanations. A typical case is illustrated in Case 3, where ChatGPT tends to prioritize the BLEU score instead of offering judgments based on its inherent capabilities.
@@ E.3 LLMs may generate invalid answers for all prompting strategies
System Online-A.en
Domain conversational
Doc_id 1
Seg_id 6
Text
Source: 请问,订单情况现在是什么样?
Reference: May I ask what the status of the order is now?
Translation: Please ask, what is the order situation now?
Human
Evaluation
Major Error: "Please ask" - Accuracy/Mistranslation
Minor Error: "situation" - Style/Awkward
Table 6: An example of MQM, comprising information of the test sample along with human-annotated major and minor errors.

### v4

@@ Abstract
Qingyu Lu,    Baopu Qiu,  Liang Ding,  Kanjian Zhang♢♠thanks:   Corresponding Author.,  Tom Kocmi,  Dacheng Tao  Southeast University  Nanjing University  The University of Sydney  Southeast University Shenzhen Research Institute  Microsoft  Nanyang Technological University [Uncaptioned image] luqingyu@seu.edu.cn, qiubaopu@smail.nju.edu.cn, liangding.liam@gmail.com, tomkocmi@microsoft.com [Uncaptioned image] https://github.com/Coldmist-Lu/ErrorAnalysis_Prompt
@@ 1 Introduction
Large language models (LLMs), especially Generative Pre-trained Transformer (GPT) models Radford et al. (2019); Brown et al. (2020) such as ChatGPT (Ouyang et al., 2022; Achiam et al., 2023), have shown remarkable performance in various natural language processing (NLP) tasks Qin et al. (2023); Zhong et al. (2023) and complex reasoning and agent applications Zhong et al. (2024); Ren et al. (2024). LLMs are capable of integrating multiple NLP tasks and can generate detailed and comprehensive responses to human inquiries. Additionally, they can respond appropriately to follow-up questions and maintain sensitivity throughout several turns of conversation.
@@ 2.1 Translation Evaluation Metric
To verify the reliability of MT metrics, Multi-dimensional Quality Metric (MQM) has been adopted recently in WMT as a high-quality human evaluation strategy (Freitag et al., 2021). It asks human experts to annotate the errors in the hypothesis and categorize them into "Major" and "Minor" indicating their severity. A detailed description of MQM annotation is presented in Appendix A.
@@ 2.3 Error Analysis Prompting
In this work, we combine the CoT and EA paradigms, introducing a novel prompting strategy called Error Analysis Prompting (EAPrompt). As shown in Figure 1, EAPrompt divides the scoring process into two stages: First, the LLM is instructed to identify major and minor errors in the translation ("Instruction: Identify Errors"). Subsequently, the number of these two types of errors is counted ("Instruction: Count Errors"). Distinguished from GEMBA prompting, EAPrompt emulates the evaluation process of MQM and produces more explainable and reliable automatic evaluations.
@@ 2.4 Post-processing of LLM responses
WMT22 En-De 2037 17 conversational, e-commerce, news, social
En-Ru 2037 17 conversational, e-commerce, news, social
Zh-En 1875 20 conversational, e-commerce, news, social
@@ 3.1 Experiment Setup
We utilize the test set from the WMT22 shared tasks (Freitag et al., 2022) in English-German (En-De), English-Russian (En-Ru), and Chinese-English (Zh-En) across 4 different domains - conversational, e-commerce, news, and social. This study necessitates the evaluation of 106,758 segments from 54 MT systems. In accordance with GEMBA Kocmi and Federmann (2023a), we deliberately exclude previous datasets such as WMT21 and WMT20. This reduces the potential risk of data contamination, as LLMs may have utilized these datasets in their training sets. Table 1 provides statistics about our test set. For future studies, we plan to incorporate more recent testsets, such as WMT23 Freitag et al. (2023).
@@
Significance Analysis
At the segment level, we follow WMT22 metrics shared task to utilize PERM-BOTH hypothesis test (Deutsch et al., 2021) to assess the significance of metrics. We use 1000 re-sampling runs and set $p=0.05$.
@@ 3.2 Baselines and Large Language Models
For prioprietary models, we use the OpenAI API to experiment with GPT-3.5-Turbo [^2] We use the 0613 OpenAI model version. Due to budget constraints, we verify EAPrompt on GPT-4 with a limited number of samples. Please refer to Appendix E.1 for more details.. We also experiment with an inhouse human-aligned Llama2-70B series model Touvron et al. (2023b) fine-tuned with multilingual translation data, noted as "Llama2-70b-Chat" in experimental results. We also use a high-quality sparse mixture-of-experts model, Mixtral-8x7b (Jiang et al., 2024). We use a state-of-the-art checkpoint Mixtral-8x7b-Instruct which has been optimised through supervised fine-tuning and direct preference optimisation to follow instructions.
@@ 3.3 Prompts for LLM evaluators
3.4 Inference Costs for EAPrompt
It is importatant to note that using a more complex prompting strategy may lead to higher inference costs. In EAPrompt, there are two steps that may require querying LLM for each translation:
Identifying Errors
This step generates a maximum of 256 tokens in total and may terminate early if all errors are listed.
Counting the Number of Errors
This step can be implemented in two ways. If querying LLM, it outputs only two numbers with a fixed output format, costing less than 10 tokens. Alternatively, if a regular expression matching algorithm is employed, this step does not require inference, thus reducing costs. As described in our experimental results later 4.3 these two approaches have minimal influence on the final performance.
@@ 3.5 Experimental Results
EAPrompt 89.4 (+2.5) 55.7†(+1.0) 53.4(+3.4) 48.8†(+1.2)
Table 2: The performance of metrics using pairwise accuracy (%) at the system level and pairwise accuracy with tie calibration (%) at the segment level. All results are compared with human-annotated MQM scores. The best results among the same model are highlighted in bold. The best results among all metrics are underlined. "†" denotes cases where one metric is significantly better than the other.
@@ 4.4 Case Study
We discuss potential issues encountered by LLMs and their corresponding solutions in Appendix F, including invalid responses, input order bias, etc. We aim to provide insights that should be considered when utilizing LLMs as translation evaluators.
@@ 5 Related Work
LLMs refers to language models with hundreds of billion of parameters which are trained on massive textual data (Chang et al., 2024; Zhao et al., 2023). Since the emergence of ChatGPT, LLMs have shown remarkable proficiency across various NLP tasks (Achiam et al., 2023; Touvron et al., 2023b). A prevalent application of LLMs is harnessing them as evaluators for assessing the performance of Chatbots (Zheng et al., 2023) or some self-criticize/ improvement procedures Zhang et al. (2024); Valmeekam et al. (2023). Recent studies also show LLM’s efficacy in evaluating NLG tasks like summarization and dialogue generation through multi-step prompting (Liu et al., 2023). GEMBA (Kocmi and Federmann, 2023b) is the pioneering effort in utilizing LLMs as translation evaluators via a zero-shot prompting approach with GPT models. In this work, EAPrompt innovatively combines the basic ideas of error analysis (Lu et al., 2023) and chain-of-thought (Wei et al., 2022) to prompt LLMs for achieving human-like translation evaluation.
@@ 6 Conclusion
In this paper, we explore the potential of LLMs as a metric for evaluating translations. We design a novel one-shot prompting strategy EAPrompt based on chain-of-thought and error analysis, and show that this strategy significantly improves the evaluation performance on both the system and segment levels. We compare different EAPrompt variants and ultimately opt for a 2-step prompting approach with itemized error demonstrations. Further analysis confirms EAPrompt’s proficiency in error identification and its alignment with the commonly accepted human evaluations MQM. In future work, we would like to experiment with a broader range of LLMs Barrault et al. (2019); Anastasopoulos et al. (2021); Kocmi et al. (2022); Zan et al. (2022), to make our conclusion more convincing. Lastly, it will be interesting to test the capabilities of LLMs for other MT-related tasks, such as grammatical error correction and automatic post-editing Wu et al. (2023); Vidal et al. (2022).
@@
Acknowledgments
We thank the anonymous reviewers and the area chair for their insightful comments and suggestions. This work was supported in part by the National Natural Science Foundation of China under Grant 61973083, and in part by the Shenzhen Science and Technology Program JCYJ20210324121213036.
@@ Appendix A Description of MQM
Multidimensional Quality Metric (MQM) is a human evaluation framework commonly used in WMT metrics shared tasks as the golden standard (Freitag et al., 2021, 2023). It is developed to evaluate and categorize errors in translations. The annotations from human experts are open-sourced and available from WMT22 metrics shared tasks for scientific research (Freitag et al., 2022).
In WMT22, MQM annotations for En-De and Zh-En were sponsored and executed by Google, using 11 professional translators (7 for En-De, 4 for Zh-En). The annotations for En-Ru were provided by Unbabel who utilized 4 professional, native language annotators with ample translation experience. They have access to the full document context.
About the inter-rater agreement. In MQM, each segment is annotated by 2 or 3 annotators. The final segment-level score is an average over scores from all annotators. As depicted in (Freitag et al., 2021), the pairwise inter-rater agreement is about 0.584 for En-De, and 0.412 for Zh-En, which is significantly better than other evaluation protocols such as Scalar Quality Metric and Direct Assessment.
In this paper, EAPrompt emulates MQM to identify major and minor errors, providing insightful explanations for the translation. Table 5 shows an example annotated through MQM framework.
System Online-A.en
Domain conversational
Doc_id 1
Seg_id 6
Source(Zh) {CJK}UTF8gbsn请问,订单情况现在是什么样?
Reference(En) May I ask what the status of the order is now?
Translation(En) Please ask, what is the order situation now?
Major Error(s) "Please ask" - Accuracy/Mistranslation
Minor Error(s) "situation" - Style/Awkward
Table 5: An example of MQM, comprising information of the test sample along with human-annotated errors.
@@ Appendix D Counting Errors using Regular Expressions Matching
Appendix E Additional Results
E.1 Results on GPT-4
In order to further verify the effect of EAPrompt on state-of-the-art LLMs, we test our approach on GPT-4 as an initial experiment, with a subset of 600 samples, due to budget constraints. Results are shown in Table 7. We can see that on powerful GPT-4, EAPrompt shows a slight performance advantage over GEMBA-DA, outperforming it by 1.05 points.
Model GEMBA EAPrompt $\delta$
GPT-4 44.92 45.97 +1.05
Table 7: Segment Level Kendall Correlation on WMT22 Zh-En using GPT-4 with a subset of 600 samples.
E.2 Results at the system-level
En-De En-Ru Zh-En
Models Prompts Ref? $\rho$ Acc. $\rho$ Acc. $\rho$ Acc.
Llama2-70b-Chat GEMBA 53.0 69.2 79.4 84.8 77.3 77.3
EAPrompt 78.3 76.9 84.2 86.7 98.2 98.2
GEMBA 61.5 53.8 83.2 84.8 93.4 74.7
EAPrompt 76.2 83.3 85.7 88.6 97.6 84.6
Mixtral-8x7b-Instruct GEMBA 35.1 59.0 64.3 73.3 89.9 74.7
EAPrompt 54.6 74.4 84.5 88.6 95.8 95.8
GEMBA 70.7 82.1 83.0 85.7 98.0 52.7
EAPrompt 71.4 78.2 83.6 86.7 97.5 82.4
GPT-3.5-Turbo GEMBA 87.2 85.9 81.7 90.5 98.4 82.4
EAPrompt 92.1 91.0 92.0 92.4 98.3 92.3
GEMBA 86.6 83.3 77.3 90.5 98.3 85.7
EAPrompt 91.9 88.5 82.4 90.5 98.2 89.0
Table 8: The performance of metrics for each language pair using pearson correlation ($\rho$) and pairwise accuracy (Acc.) at the system level. All results are compared with human-annotated MQM scores. The best results among the same model are highlighted in bold. The best results among all metrics are underlined.
To complement our main findings, we also compare our method, EAPrompt, with GEMBA using pearson correlation and pairwise accuracy. In addition to Table 2, which groups three language pairs together to present comprehensive and unbiased results, Table 8 provides a system-level performance comparison across different language pairs for a more detailed analysis.
We can see that at the system level, across all LLMs and for all three language pairs, EAPrompt consistently outperforms GEMBA in most scenarios, aligning with the primary conclusions drawn in our paper. Additionally, we note a few cases where the two metrics exhibit slight discrepancies. This discrepancy, typically less than 1.0 point, may be attributed to Pearson correlation’s sensitivity to low-quality MT systems (Mathur et al., 2020a), potentially leading to an unfair judgment compared to Pairwise Accuracy.
E.3 Results at the segment-level
En-De En-Ru Zh-En
Models Prompts Ref? $\rho$ Acc* $\rho$ Acc* $\rho$ Acc*
Llama2-70b-Chat GEMBA 10.8 53.7 6.3 48.8 2.0 45.4
EAPrompt 27.8 55.2 28.1† 51.4† 39.1 50.2
GEMBA 4.3 54.1 5.0 47.8 1.2 45.0
EAPrompt 26.2† 55.0† 29.8 51.6 38.6† 49.3†
Mixtral-8x7b-Instruct GEMBA 23.8 54.8 6.4 48.3 2.5 46.7
EAPrompt 30.1 53.8 28.5 50.6 40.2 48.2†
GEMBA 12.2 54.8 7.7 47.5 0.9 46.2
EAPrompt 26.7† 54.1 24.5† 49.9† 37.9† 48.3
GPT-3.5-Turbo GEMBA 28.6 55.2 21.0 49.5 24.8 48.2
EAPrompt 39.9 56.7 34.6 53.3† 41.9 50.0
GEMBA 29.6 54.7 24.3 50.0 24.9 47.6
EAPrompt 33.3† 55.7† 30.2† 53.4 36.2† 48.8†
Table 9: The performance of metrics for each language pair using pearson correlation ($\rho$) and pairwise accuracy with ties (Acc*) at the segment level. All results are compared with human-annotated MQM scores. The best results among the same model are highlighted in bold. The best results among all metrics are underlined. "†" denotes cases where one metric is significantly better than the other.
We present more detailed results, including Pearson correlation, as shown in Table 9. We can observe that when using Pearson correlation as the meta-evaluation method, EAPrompt surpasses GEMBA by a significant margin across all scenarios considered in our experiments. This further validates the effectiveness and strong performance of EAPrompt compared to GEMBA.
Appendix F Case Study
@@
F.1 Potential instability in the responses without temperature control
@@
F.2 Input order bias when evaluating multiple translations simultaneously
@@
F.3 LLMs may generate invalid answers for all prompting strategies
Issue: We observe that in certain cases, LLMs may not function as translation evaluators that may produce invalid answers with textual explanations. A typical case is illustrated in Case 3, where ChatGPT tends to prioritize the BLEU score instead of offering judgments based on its inherent capabilities.